OpenStudy (anonymous):

Solve the system using substitution. y=-2x 5x+y=3

OpenStudy (shamil98):

y = -2x 5x + y = 3 If y = -2x and 5x + y = 3 you just input the value for y (-2x) into that equation and simplify and solve for x. 5x +(-2x) = 3 5x - 2x = 3 and solve for x.

OpenStudy (anonymous):

1,-2

OpenStudy (shamil98):

Correct :)

OpenStudy (anonymous):

Thanks! :D